OpenStudy (anonymous):

you are choosing between two health clubs club A offers memberships for a fee of $21 plus monthly fee of $26. Club B offers a membership for a fee of $13 plus a monthly fee of $30. After how many months will the total costof each health clubbe the same?what will be the total cost for each club? In ______ months the total cost of each health club will be the same.

OpenStudy (anonymous):

please help

OpenStudy (rulnick):

21+26m = 13 + 30m 8 = 4m m = 2 The costs will be the same after 2 months ($73).